# READ: Real-time and Efficient Asynchronous Diffusion for Audio-driven Talking Head Generation

### v2

@@ Abstract
The introduction of diffusion models has brought significant advances to the field of audio-driven talking head generation. However, the extremely slow inference speed severely limits the practical implementation of diffusion-based talking head generation models. In this study, we propose READ, the first real-time diffusion-transformer-based talking head generation framework. Our approach first learns a spatiotemporal highly compressed video latent space via a temporal VAE, significantly reducing the token count to accelerate generation. To achieve better audio-visual alignment within this compressed latent space, a pre-trained Speech Autoencoder (SpeechAE) is proposed to generate temporally compressed speech latent codes corresponding to the video latent space. These latent representations are then modeled by a carefully designed Audio-to-Video Diffusion Transformer (A2V-DiT) backbone for efficient talking head synthesis. Furthermore, to ensure temporal consistency and accelerated inference in extended generation, we propose a novel asynchronous noise scheduler (ANS) for both the training and inference process of our framework. The ANS leverages asynchronous add-noise and asynchronous motion-guided generation in the latent space, ensuring consistency in generated video clips. Experimental results demonstrate that READ outperforms state-of-the-art methods by generating competitive talking head videos with significantly reduced runtime, achieving an optimal balance between quality and speed while maintaining robust metric stability in long-time generation.
@@ 1 Introduction
Audio-driven talking head generation, also known as audio-driven portrait animation, aims to generate realistic videos of a person speaking from a given audio signal, which demonstrates significant value across multiple domains such as e-learning, film and game production, and human-computer interaction (Chen et al. 2020). Evaluation criteria for audio-driven talking head generation models include the accuracy of lip synchronization with the input audio and the naturalism of the generated facial movements. In addition to these factors, the model’s inference speed is also a crucial metric, as achieving real-time capabilities is essential for future human-computer interactive applications (Zhen et al. 2023).
@@
Recently, the field of talking head generation has been greatly advanced by the introduction of diffusion models (Croitoru et al. 2023). Talking head generation frameworks built on the foundations of image or video diffusion models (Wang et al. 2025; Ji et al. 2025) achieve more vivid performance than traditional methods (Wang et al. 2021; Zhang et al. 2023b). However, existing diffusion-based talking head generation models generally suffer from extremely slow inference speed, typically requiring tens to hundreds of seconds to generate a mere 5-second video (Wang et al. 2025; Ji et al. 2025; Chen et al. 2025), presenting a new challenge to this research field. The slow inference speed can be attributed to the following factors. First, the talking head generation task necessitates temporal alignment between speech features and video latents to ensure lip-sync accuracy. Existing methods typically employ a Variational Autoencoder (VAE) (Kingma, Welling et al. 2013) without temporal compression to achieve better alignment (Ji et al. 2025; Chen et al. 2025; Xu et al. 2024), yet substantially increase the input token count and computational cost of the model. Second, conventional Denoising Diffusion Probabilistic Models (DDPM) (Nichol and Dhariwal 2021) or Denoising Diffusion Implicit Models (DDIM) (Song, Meng, and Ermon 2020) sampling methods require a large number of inference steps to generate high-fidelity video, substantially increasing inference time. Furthermore, considering extended generation, existing solutions mainly adopt overlap-and-fuse techniques (Wang et al. 2025; Ji et al. 2025) or introduce an auxiliary network (Cui et al. 2025; Xu et al. 2024) to maintain consistency between generated video clips, further increasing computational and time costs.
@@
To address this challenge, in this research we introduce READ, the first real-time diffusion-transformer-based talking head generation framework. Our framework incorporates a pre-trained temporal VAE with a high spatiotemporal compression ratio of 32×32×8 pixels per token. To achieve better audio-visual alignment in the compressed latent space, we pre-train a Speech Autoencoder (SpeechAE) by self-supervising to generate temporally compressed speech latent codes that preserve essential acoustic information corresponding to the compressed video latent space. Then, an Audio-to-Video Diffusion Transformer (A2V-DiT) is designed to generate video latents under speech latent conditions efficiently. The whole training and inference process of our framework is under the proposed Asynchronous Noise Scheduler (ANS), which implements an asynchronous add-noise forward process and an asynchronous motion-guided reverse process to effectively generate long-time videos.
@@ 2.1 Audio-driven Talking Head Generation
Audio-driven talking head generation aims to generate a talking person video conditioned on audio input, garnering increasing research interest due to its extensive application scenarios. Early research in audio-driven talking head generation primarily focused on achieving accurate lip synchronization with the input audio (Prajwal et al. 2020). Subsequent works, such as Audio2Head (Wang et al. 2021) and SadTalker (Zhang et al. 2023b), advanced the field by producing more naturalistic head movements. More recent models, including DreamTalk (Zhang et al. 2023a), Diffused Heads (Stypułkowski et al. 2024), have further enhanced the expressiveness of the generated animations. Recently, a major shift occurred with the introduction of pretrained diffusion models (Blattmann et al. 2023; Rombach et al. 2022). Frameworks like Sonic (Ji et al. 2025), EmotiveTalk (Wang et al. 2025), and Hallo (Xu et al. 2024) now leverage these powerful image or video diffusion priors to generate videos with improved fidelity and realism. However, a critical limitation of these models is their slow inference speed. We address this issue by proposing a novel framework specifically designed for fast talking head generation.
@@ 2.2 Fast Diffusion Models
Accelerating diffusion models is a major research focus, with two dominant strategies: reducing inference steps and optimizing the model architecture (Shen et al. 2025). The first strategy includes methods such as Progressive Distillation (Salimans and Ho 2022), ADD (Sauer et al. 2024b), LADD (Sauer et al. 2024a) and others (Meng et al. 2023; Yin et al. 2024). The second strategy involves lightweight networks or VAEs with high compression ratios. In particular, video generation methods such as LTX-VIDEO (HaCohen et al. 2024) and Wan (Wan et al. 2025) have successfully used temporal and spatial compression in their VAEs to reduce computational cost. However, a critical issue arises when applying these architectural optimizations to talking head generation, as temporal compression undermines the audio-video alignment essential for accurate lip synchronization. To address this, we introduce a SpeechAE with self-supervised pre-training for synchronous speech feature compression to achieve better audio-visual alignment within latent space, and an Asynchronous Noise Scheduler (ANS) designed to ensure fast and stable extended inference.
@@ 3.1 Preliminary
Flow Matching. Define ${\mathit{\boldsymbol{Z}}}(0)$ as the original latents obtained by VAE, and ${\mathit{\boldsymbol{Z}}}(t)$ as the noisy latents at timestep $t$. Flow Matching (FM) (Lipman et al. 2022) is a generative method that leverages the principles of Ordinary Differential Equations (ODEs) (Hartman 2002). The central idea is to learn a continuous-time vector field ${\mathit{\boldsymbol{v}}}({\mathit{\boldsymbol{Z}}}(t),t)$ that transports samples from a simple noise distribution to the target data distribution ${\mathit{\boldsymbol{Z}}}(0)$ (Lipman et al. 2024; Dao et al. 2023):
@@ 3.2 Fast Audio-to-Video Generation Framework
Temporal VAE for Video Compression. Training and inference time for DiT models is dominated by the number of input tokens (Peebles and Xie 2023). To reduce the number of tokens processed by the backbone network to accelerate generation speed, we employ a temporal VAE with a high spatiotemporal compression ratio of 32×32×8 pixels per token from LTX-VIDEO (HaCohen et al. 2024). The principle can be formulated as follows:
@@
SpeechAE for Speech Feature Compression. Unlike the text-to-video generation task, achieving precise temporal alignment between speech and video latents is particularly critical to achieve accurate lip synchronization in talking head generation. Although temporal VAE achieves a high compression ratio, it hinders audio-visual alignment because temporal compression disrupts the original correspondence between video and speech signals. We proposed SpeechAE self-supervised pre-training to perform synchronous temporal compression on raw speech features to address this limitation. Shown in (b) of Fig. 1. Our SpeechAE integrates a frozen Whisper-tiny encoder (Radford et al. 2023) for speech feature extraction, as described below:
@@
The trainable part of SpeechAE also employs an encoder-decoder architecture, which consists of linear-based dimensionality transformation modules and temporal sampling modules built with 1D causal convolutional (Li et al. 2021) layers, achieving the same temporal compression ratio as $\mathcal{E}_{\text{V}}$. The compressed speech latent codes $\mathit{\boldsymbol{C}\in\mathbb{R}^{f\times h_{w}\times d_{\text{A}}}}$ are generated by the SpeechAE encoder from $\mathit{\boldsymbol{S}}=[\mathit{\boldsymbol{s}}_{1},...,\mathit{\boldsymbol{s}}_{F}]\in\mathbb{R}^{F\times H_{w}\times D_{\text{A}}}$, and reconstructed to $\hat{\mathit{\boldsymbol{S}}}=[\hat{\mathit{\boldsymbol{s}}}_{1},...,\hat{\mathit{\boldsymbol{s}}}_{F}]$ through the decoder, as follows:
@@
where $H_{w}$ and $h_{w}$ indicates the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denotes the hidden dims. The quality of reconstruction can serve as a proxy for the information lost during compression (Wang, Yao, and Zhao 2016). Effective reconstruction indicates that the latent codes $\mathit{\boldsymbol{C}}$ retain critical temporal information of source features $\mathit{\boldsymbol{S}}$}. Based on this, we introduce a self-supervised pre-training phase on SpeechAE for the task of auto-encoding. To minimize the Euclidean distance between raw features $\mathit{\boldsymbol{S}}$} and reconstructed features $\hat{\mathit{\boldsymbol{S}}}$, we first apply a Mean Squared Error (MSE) loss, as follows:
@@ 3.3 Asynchronous Noise Scheduler (ANS)
Then we sample the asynchronous noise timestep $\boldsymbol{t}$ from the shifted-logit-normal distribution (Esser et al. 2024) based on the aforementioned latent structure that applies different noise timesteps to motion and reference frames, as follows:
@@
The generated optical flow is then mapped back to the latent space via the FM scheduler (Lipman et al. 2022). This process iterates until all timesteps are processed, resulting in the generated temporally consistent latent sequence $\mathit{\boldsymbol{Z}}(0)$.
@@ 4.1 Experimental Setup
Implementation Details. Experiments encompassing both training and inference are conducted on the open-source HDTF (Zhang et al. 2021) and MEAD (Wang et al. 2020) datasets. 95% data of both datasets is randomly allocated for training and the remaining 5% for testing, ensuring no overlap between the partitions. We employ a two-stage training strategy. In the first stage, the SpeechAE module is pre-trained with a learning rate of $1\times 10^{-4}$. In the second stage, the entire audio-to-video backbone is trained at a resolution of $512\times 512$ pixels and 121 frames, with a learning rate of $1\times 10^{-5}$ and a batch size of 1. All reported results use 8-step sampling and Split-CFG with $\alpha=2.0$ and $\beta=6.0$ unless specified, and enable step reduction down to 5 steps due to the efficiency of the proposed ANS scheduler. The inference window length is set to match the training length with a motion overlap of one frame in latent space. Both training and evaluation are performed on NVIDIA A100 GPUs. Other details about the dataset and the training and inference of our framework can be found in the supplementary PDF.
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (Seitzer 2020) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (Unterthiner et al. 2019) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (Chung and Zisserman 2017), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (Tian et al. 2024) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
Baselines. We benchmark our proposed method against several state-of-the-art (SOTA) open-source implementations, including full-video diffusion strategies such as Sonic (Ji et al. 2025), EchoMimic (Chen et al. 2025), Hallo (Xu et al. 2024), and AniPortrait (Wei, Yang, and Wang 2024), as well as motion-space diffusion models like AniTalker (Liu et al. 2024). All quantitative and qualitative comparisons are conducted on the same device using identical test data with the same length of 4.84s (121 frames) to ensure fair evaluation.
@@ 4.2 Overall Evaluation
As demonstrated in Table 1, the representative motion-space diffusion method AniTalker achieves reduced runtime compared to other full-video diffusion approaches, yet underperforms full-video diffusion methods on FID and FVD due to the absence of a diffusion-based renderer. In contrast, our full-video diffusion approach achieves substantially lower latency in the backbone runtime than other methods, which represents a significant step forward for the acceleration of diffusion-based talking head generation. In addition to its speed, our approach also achieves highly competitive performance across all the evaluation metrics. On both HDTF and MEAD datasets, our model surpasses all competing methods in terms of FID and Sync-C, while it also achieves SOTA or near-SOTA performance on FVD, Sync-D, and E-FID. Notably, our model establishes superior E-FID performance on the emotion-rich MEAD dataset, validating its capability for generating expressive expressions faithful to the speech.

### v3

@@ Abstract
The introduction of diffusion models has brought significant advances to the field of audio-driven talking head generation. However, the extremely slow inference speed severely limits the practical implementation of diffusion-based talking head generation models. In this study, we propose READ, a real-time diffusion-transformer-based talking head generation framework. Our approach first learns a spatiotemporal highly compressed video latent space via a temporal VAE, significantly reducing the token count to accelerate generation. To achieve better audio-visual alignment within this compressed latent space, a pre-trained Speech Autoencoder (SpeechAE) is proposed to generate temporally compressed speech latent codes corresponding to the video latent space. These latent representations are then modeled by a carefully designed Audio-to-Video Diffusion Transformer (A2V-DiT) backbone for efficient talking head synthesis. Furthermore, to ensure temporal consistency and accelerated inference in extended generation, we propose a novel asynchronous noise scheduler (ANS) for both the training and inference processes of our framework. The ANS leverages asynchronous add-noise and asynchronous motion-guided generation in the latent space, ensuring consistency in generated video clips. Experimental results demonstrate that READ outperforms state-of-the-art methods by generating competitive talking head videos with significantly reduced runtime, achieving an optimal balance between quality and speed while maintaining robust metric stability in long-time generation.
Project Pagehttps://readportrait.github.io/READ
@@ 1 Introduction
Audio-driven talking head generation aims to generate videos of a person speaking an audio signal, which demonstrates significant value across multiple domains such as e-learning, film and game production, and human-computer interaction (thdsurvey1). Evaluation criteria for audio-driven talking head generation models include the accuracy of lip synchronization with the input audio and the naturalism of the generated facial movements. In addition to these factors, the model’s inference speed is also a crucial metric, as achieving real-time capabilities is essential for future human-computer interactive applications (thdsurvey2).
@@
Recently, the field of talking head generation has been greatly advanced by the introduction of diffusion models (diffusionsurvey). Talking head generation frameworks built on the foundations of image or video diffusion models (emotivetalk) achieve more vivid performance than traditional methods (audio2head; sadtalker). However, existing diffusion-based talking head generation models generally suffer from extremely slow inference speed, typically requiring tens to hundreds of seconds to generate a mere 5-second video (sonic; echomimic), presenting a new challenge to this research field. The slow inference speed can be attributed to the following factors. First, the talking head generation task necessitates temporal alignment between speech features and video latents to ensure lip-sync accuracy. Existing methods typically employ a Variational Autoencoder (VAE) (vae) without temporal compression to achieve better alignment (sonic; echomimic; hallo), yet increase the input token count and computational cost of the model. Second, conventional Denoising Diffusion Probabilistic Models (DDPM) (ddpm) or Denoising Diffusion Implicit Models (DDIM) (ddim) sampling methods require a large number of inference steps to generate high-fidelity video, substantially increasing inference time. Furthermore, considering extended generation, existing solutions mainly adopt overlap-and-fuse techniques (emotivetalk; sonic) or introduce an auxiliary network (hallo3; hallo) to maintain consistency between generated video clips, further increasing computational and time costs.
@@
To address this challenge, in this research we introduce READ, the first end-to-end real-time diffusion-transformer-based audio-driven talking head generation framework. Our framework incorporates a temporal VAE with a high compression ratio of 32×32×8 pixels per token. To achieve better audio-visual alignment in the compressed latent space, we pre-train a Speech Autoencoder (SpeechAE) by self-supervising to generate temporally compressed speech latent codes corresponding to the compressed video latents. Then, an Audio-to-Video Diffusion Transformer (A2V-DiT) is designed to generate video latents under speech latent conditions efficiently. The training and inference processes of our framework are under the proposed Asynchronous Noise Scheduler (ANS), which implements an asynchronous add-noise forward process and an asynchronous motion-guided reverse process to effectively generate long-time videos.
@@ 2.1 Audio-driven Talking Head Generation
Audio-driven talking head generation aims to generate a talking person video conditioned on audio input, garnering increasing research interest due to its extensive application scenarios. Early research in audio-driven talking head generation primarily focused on achieving accurate lip synchronization with the input audio (wav2lip). Subsequent works, such as Audio2Head (audio2head) and SadTalker (sadtalker), advanced the field by producing more naturalistic head movements. More recent models, including DreamTalk (dreamtalk), Diffused Heads (diffusedheads), have further enhanced the expressiveness of the generated animations. Recently, a major shift occurred with the introduction of pretrained diffusion models (stablevideodiffusion; latentdiffusion). Frameworks like Sonic (sonic), EmotiveTalk (emotivetalk), and Hallo (hallo) now leverage these powerful image or video diffusion priors to generate videos with improved fidelity and realism. However, a critical limitation of these models is their slow inference speed. We address this issue by proposing a novel framework specifically designed for fast talking head generation.
@@ 2.2 Fast Diffusion Models
Accelerating diffusion models is a major research focus (accelerationsurvey). Progressive Distillation (progressivedistillation), ADD (add), LADD (ladd) and others (guideddiffusion; onestepdiffusion) focus on reducing diffusion steps. Ditto (ditto) and AniTalker (anitalker) employ motion-space diffusion to reduce tokens processed by the diffusion backbone for acceleration, yet face challenges with the naturalness of the generated video. In contrast, end-to-end video generation methods such as LTX-VIDEO (ltxvideo) and Wan (wan) utilize spatiotemporal compression in their VAEs to reduce computational cost. However, a critical issue arises when applying these VAEs to talking head generation, as temporal compression undermines the audio-visual alignment essential for accurate lip synchronization. To address this, we introduce a SpeechAE with self-supervised pre-training for synchronous speech feature compression to achieve better audio-visual alignment within end-to-end diffusion, and an Asynchronous Noise Scheduler (ANS) designed to ensure fast and stable extended inference.
@@ 3.1 Preliminary
Flow Matching. Define ${\mathit{\boldsymbol{Z}}}(0)$ as the original latents obtained by VAE, and ${\mathit{\boldsymbol{Z}}}(t)$ as the noisy latents at timestep $t$. Flow Matching (FM) (flowmatching) is a generative method that leverages the principles of Ordinary Differential Equations (ODEs) (ode). The central idea is to learn a continuous-time vector field ${\mathit{\boldsymbol{v}}}({\mathit{\boldsymbol{Z}}}(t),t)$ that transports samples from a simple noise distribution to the target data distribution ${\mathit{\boldsymbol{Z}}}(0)$ (flowmatching1; flowmatching2):
@@ 3.2 Fast Audio-to-Video Generation Framework
Temporal VAE for Video Compression. Training and inference time for DiT models is dominated by the number of input tokens (dit). To reduce the number of tokens processed by the backbone network to accelerate generation speed, we employ a temporal VAE with a high spatiotemporal compression ratio of 32×32×8 pixels per token from LTX-VIDEO (ltxvideo). The principle can be formulated as follows:
@@
SpeechAE for Speech Feature Compression. Unlike the text-to-video generation task, achieving precise temporal alignment between speech and video latents is particularly critical to achieve accurate lip synchronization in talking head generation. Although temporal VAE achieves a high compression ratio, it hinders audio-visual alignment because temporal compression disrupts the original correspondence between video and speech signals. We proposed SpeechAE self-supervised pre-training to perform synchronous temporal compression on raw speech features to address this limitation. Shown in (b) of Fig. 1. Our SpeechAE integrates a frozen Whisper-tiny encoder (whisper) for speech feature extraction, as described below:
@@
The trainable part of SpeechAE also employs an encoder-decoder architecture, which consists of linear-based dimensionality transformation modules and temporal sampling modules built with 1D causal convolutional (cnn) layers, achieving the same temporal compression ratio as $\mathcal{E}_{\text{V}}$. The compressed speech latent codes $\mathit{\boldsymbol{C}\in\mathbb{R}^{f\times h_{w}\times d_{\text{A}}}}$ are generated by the SpeechAE encoder from $\mathit{\boldsymbol{S}}=[\mathit{\boldsymbol{s}}_{1},...,\mathit{\boldsymbol{s}}_{F}]\in\mathbb{R}^{F\times H_{w}\times D_{\text{A}}}$, and reconstructed to $\hat{\mathit{\boldsymbol{S}}}=[\hat{\mathit{\boldsymbol{s}}}_{1},...,\hat{\mathit{\boldsymbol{s}}}_{F}]$ through the decoder, as follows:
@@
where $H_{w}$ and $h_{w}$ indicates the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denotes the hidden dims. The quality of reconstruction can serve as a proxy for the information lost during compression (ae). Effective reconstruction indicates that the latent codes $\mathit{\boldsymbol{C}}$ retain critical temporal information of source features $\mathit{\boldsymbol{S}}$}. Based on this, we introduce a self-supervised pre-training phase on SpeechAE for the task of auto-encoding. To minimize the Euclidean distance between raw features $\mathit{\boldsymbol{S}}$} and reconstructed features $\hat{\mathit{\boldsymbol{S}}}$, we first apply a Mean Squared Error (MSE) loss, as follows:
@@ 3.3 Asynchronous Noise Scheduler (ANS)
Then we sample the asynchronous noise timestep $\boldsymbol{t}$ from the shifted-logit-normal distribution (sd3) based on the aforementioned latent structure that applies different noise timesteps to motion and reference frames, as follows:
@@
The generated optical flow is then mapped back to the latent space via the FM scheduler (flowmatching). This process iterates until all timesteps are processed, resulting in the generated temporally consistent latent sequence $\mathit{\boldsymbol{Z}}(0)$.
@@ 4.1 Experimental Setup
Implementation Details. Experiments encompassing both training and inference are conducted on HDTF (hdtf) and MEAD (mead) datasets. 95% data of both datasets is randomly allocated for training and the remaining 5% for testing, ensuring no overlap between the partitions. We employ a two-stage training strategy. In the first stage, the SpeechAE is pre-trained with a learning rate of $1\times 10^{-4}$. In the second stage, the entire audio-to-video backbone is trained at a resolution of $512\times 512$ pixels and 121 frames, with a learning rate of $1\times 10^{-5}$ and a batch size of 1. All reported results use 8-step sampling and Split-CFG with $\alpha=2.0$ and $\beta=6.0$ unless specified. The inference window length is set to match the training length with a motion overlap of one frame in latent space. Both training and evaluation are performed on NVIDIA A100 GPUs.
Evaluation Metrics. Generation performance is assessed using several metrics. For visual quality, we employ the Fréchet Inception Distance (FID) (fid) for image-level realism between synthesized videos and reference images and the Fréchet Video Distance (FVD) (fvd) for frame-level realism between synthesized and ground-truth videos; lower values indicate better performance for both metrics. Lip synchronization is measured with SyncNet (syncnet), where a higher Synchronization Confidence (Sync-C) and a lower Synchronization Distance (Sync-D) indicate superior alignment with speech input. We further use the Expression-FID (E-FID) metric from EMO (emo) to measure the expression divergence between synthesized and ground-truth videos, with lower values indicating more faithful reproduction of expressions. Finally, we evaluate the efficiency of the diffusion backbone of each model by measuring the average runtime of the backbone per video (Runtime).
Baselines. We benchmark our method against several SOTA open-source methods, including end-to-end diffusion methods such as Sonic (sonic), EchoMimic (echomimic), Hallo (hallo), FantasyTalking (fantasytalking) and AniPortrait (aniportrait), as well as motion-space diffusion methods like AniTalker (anitalker) and Ditto (ditto). All comparisons are conducted on the same device using identical test data with the same length of 4.84s (121 frames) to ensure fair evaluation.
@@ 4.2 Overall Evaluation
As demonstrated in Tab. 1, motion-space diffusion methods such as AniTalker and Ditto achieve reduced runtime compared to other end-to-end diffusion approaches. In contrast, our end-to-end approach achieves substantially lower latency in the backbone runtime than other methods, which represents a significant step forward for the acceleration of diffusion-based talking head generation. In addition to its speed, our approach also achieves highly competitive performance across all the evaluation metrics. On both HDTF and MEAD datasets, our model surpasses all competing methods in terms of Sync-C, while it also achieves SOTA or near-SOTA performance on FID, FVD, Sync-D, and E-FID. Notably, our model establishes superior E-FID performance on the emotion-rich MEAD dataset, validating its capability for generating expressive expressions faithful to the speech.